OpenStudy (anonymous):

How do I get the derivative of this function? C=te^-4t.

OpenStudy (anonymous):

Are you familiar with the product rule for differentiation?

OpenStudy (anonymous):

ummm......d/dx(t) + d/dx(e^-4t) ?

OpenStudy (anonymous):

simply derivative of e^x is same if it is w.r.t x use product rule and the answer is e^-4t(1-4t) :)

OpenStudy (anonymous):

You have function C = f(t) = g(t)h(t). C' = f'(t) = g(t)h'(t) + g'(t)h(t)

OpenStudy (anonymous):

With g(t) = t and h(t) = e^(-4t)

OpenStudy (anonymous):

Ok what if C=Concentration and t=hours and I'm trying to find the change in concentration at time t

OpenStudy (anonymous):

You're on the right track, but it starts with the product rule and the steps I outlined. Once you have the general derivative, you can get a specific value at a specific value for "t".

OpenStudy (anonymous):

ok, so after i do the product rule amd I trying to solve for t? What happens to the c?

OpenStudy (anonymous):

With the derivative, C becomes C' which is f ' (t)

OpenStudy (anonymous):

c'=t(e^-4t)(-4) + 1(e^-4t) ? then.... c'=-4t(e^-4t) ?

OpenStudy (anonymous):

ln 4t=ln e^-4t ln4t=-4t ln4t/-4=t or am I jus trying to find an equation for the rate of change?

OpenStudy (anonymous):

c'=(1 - 4t)(e^-4t) You were very close. Now, you can use any particular "t".

OpenStudy (anonymous):

I don't have a "t" it just says to find rate of change of concentration at time t

OpenStudy (anonymous):

You just said " ... at time t" So, you either have a general "t" or a specific value for "t"

OpenStudy (anonymous):

just a general "t" so would I leave it as c'=(1-4t)(e^-4t) ?

OpenStudy (anonymous):

That's it! Sometimes no specific value is given, in which case you go with the general solution.

OpenStudy (anonymous):

ok phew! Thank you so much!

OpenStudy (anonymous):

Nice working with you. And good luck in all of your studies!